# A Framework for Objective-Driven Dynamical Stochastic Fields

### v1

@@ Abstract
Fields offer a versatile approach for describing complex systems composed of interacting and dynamic components. In particular, some of these dynamical and stochastic systems may exhibit goal-directed behaviors aimed at achieving specific objectives, which we refer to as intelligent fields. However, due to their inherent complexity, it remains challenging to develop a formal theoretical description of such systems and to effectively translate these descriptions into practical applications. In this paper, we propose three fundamental principles—complete configuration, locality, and purposefulness—to establish a theoretical framework for understanding intelligent fields. Moreover, we explore methodologies for designing such fields from the perspective of artificial intelligence applications. This initial investigation aims to lay the groundwork for future theoretical developments and practical advances in understanding and harnessing the potential of such objective-driven dynamical stochastic fields.
@@ 1 Introduction
A field refers to a theoretical framework that assigns configurations to every point in spacetime, capturing local interactions. Fields are ubiquitous across a wide range of complex systems, where intricate global behaviors can emerge from simple local rules. In particular, we are interested in dynamical and stochastic fields that evolve to achieve specific objectives. We refer to such fields as intelligent fields, reflecting their capacity to model objective-driven behavior in systems ranging from artificial intelligence to neural processes.
@@
Our approach begins with postulating three principles that could underpin such a system. The first principle asserts that there is an evolving configuration that completely characterizes the dynamics of the system. The second principle, locality, suggests that the dynamics occur within a spatial-temporal context. Lastly, the principle of purposefulness proposes that the behavior of the system is directed by objectives. Although seemingly simple, these principles lead us to develop a framework for objective-driven dynamical stochastic fields, i.e., the intelligent fields. We make formal mathematical development of the framework and discuss methodologies for designing the behavior of these fields with a view towards potential applications in artificial intelligence.
@@
The intuition for the first principle relates to the concept of a complete configuration. Consider a system being described by a time-varying configuration. If the configuration is complete, meaning that it does not require additional information to determine the system, then the current configuration is sufficient to describe the future. We consider stochastic systems under realism, meaning that these configurations evolve probabilistically and exist independently of observation. Consequently, a complete configuration at any given moment must fully determine the future probabilistic evolution of the system without relying on past configurations, making the dynamics Markovian. In particular, the configuration being complete implies that the transition rule governing the evolution of the configuration can be formulated to be time-invariant, i.e., homogeneous. Otherwise, additional external information would be necessary to define how the transition rules themselves change over time, contradicting the assumption of completeness. The following statement summarizes the above intuition.
@@
It would be quite boring if it were merely a large, plain configuration evolving on its own. Instead, imagine the system as composed of two entities, each characterized by its own configuration. Each entity’s configuration has two parts: one private and one shared. The private part is accessible only to the entity itself, whereas the shared part is observable by both entities. Thus, at any given moment, each entity observes its own private configuration along with the shared configurations from both entities, but not the other’s private configuration. As a result, each entity evolves based solely on this partial information.
@@
Following this intuition, we call this value the objective value, as a generalization of the aforementioned concepts, and demand our field to satisfy the minimization of this objective value.
@@
In this paper, we present a formal theoretical framework for analyzing and understanding such a system, which we term an intelligent field, reflecting its capacity to model objective-driven behavior. This framework will become concrete in the following sections, as we make specific choices to mathematically define the three principles: complete configuration, locality, and purposefulness.
@@
  • Subsequently, in Section 3, we detail the purposefulness principle, completing the dynamical stochastic field with objective-driven behaviors, and explore methodologies for designing the behavior of the system.
@@ 2 Dynamical Stochastic Fields
Let us build up the theoretical tools for the dynamical stochastic fields, derived from the first two principles: complete configuration and locality. To begin, we need to give a concrete definition of the first principle.
@@
Notation. Vectors and operators are represented in bold font. Additional notation will be introduced throughout the paper as needed.
@@
$\{{\hat{\bm{e}}}_{{\omega}}\}_{{\omega}\in{\Omega}},\qquad\forall{\omega},{\omega}^{\prime}\in{\Omega}:\langle{\hat{\bm{e}}}_{{\omega}},{\hat{\bm{e}}}_{{\omega}^{\prime}}\rangle=\delta_{{\omega}}^{{\omega}^{\prime}},$ (2.1)
@@
${\mathcal{H}}({\Omega})=\texttt{span}\left(\{{\hat{\bm{e}}}_{{\omega}}\}_{{\omega}\in{\Omega}}\right).$ (2.2)
Moreover, the basis set $\{{\hat{\bm{e}}}_{{\omega}}\}_{{\omega}\in{\Omega}}$ also gives a basis set for the dual space of ${\mathcal{H}}({\Omega})$, i.e., the space ${\mathcal{H}}^{*}({\Omega})$ of continuous linear functionals mapping ${\mathcal{H}}({\Omega})\to{\mathbb{R}}$, and the dual basis set is
$\{{{\check{\bm{\eta}}}}^{\omega}\}_{{\omega}\in{\Omega}},\qquad\text{where }{{\check{\bm{\eta}}}}^{\omega}:{\mathcal{H}}({\Omega})\to{\mathbb{R}}\text{ is defined by }{{\check{\bm{\eta}}}}^{{\omega}}({\hat{\bm{\varphi}}})=\langle{\hat{\bm{e}}}_{\omega},{\hat{\bm{\varphi}}}\rangle.$ (2.3)
An element ${\hat{\bm{\varphi}}}\in{\mathcal{H}}({\Omega})$ living in the Hilbert space is denoted by the symbol $\hat{\cdot}$ to distinguish it from other variables. Similarly, we use the symbol ${\check{\cdot}}$ to emphasize that the element lives in the dual space ${\mathcal{H}}^{*}({\Omega})$.
@@
The dynamics of the system can be described by dynamics within this Hilbert space. Concretely, elements ${\hat{\bm{\varphi}}}\in{\mathcal{H}}$ can be used to represent a probabilistic distribution over ${\Omega}$, as we will demonstrate after introducing some notation.
Notation. We use the Einstein notation, where any index that appears in both the superscript and subscript of a product is implicitly summed over. For example, ${\hat{\bm{\varphi}}}\in{\mathcal{H}}$ is represented by coefficients ${\varphi}^{{\omega}}\in{\mathbb{R}}$, as:
${\hat{\bm{\varphi}}}={\varphi}^{{\omega}}{\hat{\bm{e}}}_{{\omega}}:=\sum_{{\omega}\in{\Omega}}{\varphi}^{{\omega}}{\hat{\bm{e}}}_{{\omega}}.$ (2.4)
@@
A vector ${\hat{\bm{\varphi}}}\in{\mathcal{H}}$ is normalized if it represents a probability distribution, i.e.,
${\varphi}^{{\omega}}\geq 0\quad\text{and}\quad\sum_{{\omega}\in{\Omega}}{\varphi}^{{\omega}}=1.$ (2.5)
Thus, given a normalized vector ${\hat{\bm{\varphi}}}$ that describes the probability distribution of a system at a particular time, the probability of finding the system in configuration ${\omega}$ is
${\varphi}^{{\omega}}=\langle{\hat{\bm{e}}}_{{\omega}},{\hat{\bm{\varphi}}}\rangle\in[0,1].$ (2.6)
@@
Notably, as ${\Delta t}\to 0$, the above quantity converges to the delta function $\delta_{{\omega}}^{{\omega}^{\prime}}$. Since the evolution of the system is Markovian, the process can be characterized by its infinitesimal generator. To be self-contained, we define this operator as follows and show that it governs the evolution of the system.
@@
${{\mathbf{G}}}={\hat{\bm{e}}}_{{\omega}^{\prime}}G^{{\omega}^{\prime}}_{{\omega}}{{\check{\bm{\eta}}}}^{{\omega}},$ (2.8)
where $G^{{\omega}^{\prime}}_{{\omega}}\in{\mathbb{R}}$, and that repeated indices are automatically summed over. This means that a linear operator ${{\mathbf{G}}}$ is determined by
$G^{{\omega}^{\prime}}_{{\omega}}=\langle{\hat{\bm{e}}}_{{\omega}^{\prime}},{{\mathbf{G}}}{\hat{\bm{e}}}_{{\omega}}\rangle.$ (2.9)
@@
$G^{{\omega}^{\prime}}_{{\omega}}=\langle{\hat{\bm{e}}}_{{\omega}^{\prime}},{{\mathbf{G}}}{\hat{\bm{e}}}_{{\omega}}\rangle=\lim_{{\Delta t}\to 0}\frac{p_{{\Delta t}}({\omega}^{\prime}\mid{\omega})-\delta_{{\omega}}^{{\omega}^{\prime}}}{{\Delta t}}.$ (2.10)
@@
The above infinitesimal generator encapsulates all the information we need to characterize the dynamics of ${\hat{\bm{\varphi}}}(t)$, which describes the distribution of system configurations at time $t$. It provides a highly compact formulation of the system dynamics, as shown below. In fact, it is also a reformulation of the Kolmogorov forward equations (Kolmogoroff, 1931; Feller, 1940).
@@
$\frac{\mathop{}\!\mathrm{d}}{{\mathop{}\!\mathrm{d}t}}{\hat{\bm{\varphi}}}(t)={{\mathbf{G}}}{\hat{\bm{\varphi}}}(t).$ (2.11)
@@
${\hat{\bm{\varphi}}}(t)=e^{{{\mathbf{G}}}t}{\hat{\bm{\varphi}}}(0).$ (2.12)
Note that the exponential of the linear operator is defined as $e^{{{\mathbf{G}}}t}=1+\sum_{n=1}^{\infty}\frac{1}{n!}({{\mathbf{G}}}t)^{n}$. When the context is clear, we may simply use ${\hat{\bm{\varphi}}}$ to denote ${\hat{\bm{\varphi}}}(t)$ for notational ease.
Therefore, we have two formulations of the system dynamics: the dynamics of the configuration ${\omega}(t)\in{\Omega}$ itself and the dynamics of the probability distribution ${\hat{\bm{\varphi}}}(t)\in{\mathcal{H}}$. To distinguish the two formulations, we adopt the following convention.
Notation. We refer to ${\omega}(t)\in{\Omega}$ as a configuration of the system dynamics and a normalized vector ${\hat{\bm{\varphi}}}(t)\in{\mathcal{H}}$ as a state of the system dynamics. A state represents a probability distribution of configurations.
@@ 2.1 A Two-entity View
The locality principle implies that, at time $t$, entity $x$’s immediate next behavior $(\alpha_{t+{\Delta t}},\mu_{t+{\Delta t}})$ only depends on the current observed local configurations $(\alpha_{t},\beta_{t},\mu_{t})$, and entity $y$’s next behavior $(\beta_{t+{\Delta t}},\nu_{t+{\Delta t}})$ only depends on the current observed local configurations $(\alpha_{t},\beta_{t},\nu_{t})$.
Formally, the locality principle implies the immediate independence relation as shown below. Following the notation as stated in equation 2.7, we have
@@
$\langle{\hat{\bm{e}}}_{\alpha^{\prime}\mu^{\prime}},{{\mathbf{M}}}{\hat{\bm{e}}}_{\alpha\beta\mu}\rangle=\lim_{{\Delta t}\to 0}\frac{p_{{\Delta t}}(\alpha^{\prime}\mu^{\prime}\mid\alpha\beta\mu)-\delta_{\alpha\mu}^{\alpha^{\prime}\mu^{\prime}}}{{\Delta t}}.$ (2.17)
@@
$\langle{\hat{\bm{e}}}_{\beta^{\prime}\nu^{\prime}},{{\mathbf{N}}}{\hat{\bm{e}}}_{\alpha\beta\nu}\rangle=\lim_{{\Delta t}\to 0}\frac{p_{{\Delta t}}(\beta^{\prime}\nu^{\prime}\mid\alpha\beta\nu)-\delta_{\beta\nu}^{\beta^{\prime}\nu^{\prime}}}{{\Delta t}}.$ (2.18)
@@
Notation. We adopt the convention that all tensor products are automatically ordered, e.g., ${\mathcal{H}}({\mathcal{A}})\otimes{\mathcal{H}}({\mathcal{M}})\otimes{\mathcal{H}}({\mathcal{B}})\otimes{\mathcal{H}}({\mathcal{N}})$ is equivalent to ${\mathcal{H}}({\mathcal{A}})\otimes{\mathcal{H}}({\mathcal{B}})\otimes{\mathcal{H}}({\mathcal{M}})\otimes{\mathcal{H}}({\mathcal{N}})$. Therefore, the tensor product of basis vectors ${\hat{\bm{e}}}_{\alpha\mu}\in{\mathcal{H}}({\mathcal{A}}\times{\mathcal{M}})$ and ${\hat{\bm{e}}}_{\beta\nu}\in{\mathcal{H}}({\mathcal{B}}\times{\mathcal{N}})$ is ${\hat{\bm{e}}}_{\alpha\mu}\otimes{\hat{\bm{e}}}_{\beta\nu}={\hat{\bm{e}}}_{\alpha\beta\mu\nu}\in{\mathcal{H}}$. Formally, this means we work in a quotient space obtained by identifying tensor products differing only by ordering. For ease of notation, however, we omit explicit references to this quotient operation and simply adopt this ordering as a standard convention throughout this paper.
@@
Given infinitesimal generator ${{\mathbf{M}}}\colon{\mathcal{H}}({\mathcal{A}}\times{\mathcal{B}}\times{\mathcal{M}})\to{\mathcal{H}}({\mathcal{A}}\times{\mathcal{M}})$ of the entity $x$, and ${{\mathbf{N}}}\colon{\mathcal{H}}({\mathcal{A}}\times{\mathcal{B}}\times{\mathcal{N}})\to{\mathcal{H}}({\mathcal{B}}\times{\mathcal{N}})$ of the entity $y$, we define their embeddings $\bar{{\mathbf{M}}},\bar{{\mathbf{N}}}$ in the space ${\mathcal{B}}({\mathcal{H}},{\mathcal{H}})$ as the following. Given basis vectors ${\hat{\bm{e}}}_{\alpha\beta\mu\nu}\in{\mathcal{H}}$,
$\bar{{\mathbf{M}}}{\hat{\bm{e}}}_{\alpha\beta\mu\nu}$ $:=({{\mathbf{M}}}{\hat{\bm{e}}}_{\alpha\beta\mu})\otimes{\hat{\bm{e}}}_{\beta\nu},$ (2.20)
$\bar{{\mathbf{N}}}{\hat{\bm{e}}}_{\alpha\beta\mu\nu}$ $:=({{\mathbf{N}}}{\hat{\bm{e}}}_{\alpha\beta\nu})\otimes{\hat{\bm{e}}}_{\alpha\mu}.$ (2.21)
@@
We have so far considered the case of a system consisting of two entities. Next, we extend this theoretical model to multiple interacting entities, which collectively form a field.

### v2

@@ Abstract
Fields offer a versatile approach for describing complex systems composed of interacting and dynamic components. In particular, some of these dynamical and stochastic systems may exhibit goal-directed behaviors aimed at achieving specific objectives, which we refer to as intelligent fields. However, due to their inherent complexity, it remains challenging to develop a formal theoretical description of such systems and to effectively translate these descriptions into practical applications. In this paper, we propose three fundamental principles to establish a theoretical framework for understanding intelligent fields: complete configuration, locality, and purposefulness. Moreover, we explore methodologies for designing such fields from the perspective of artificial intelligence applications. This initial investigation aims to lay the groundwork for future theoretical developments and practical advances in understanding and harnessing the potential of such objective-driven dynamical stochastic fields.
@@ 1 Introduction
A field refers to a theoretical framework that assigns configurations to every point in spacetime, capturing local interactions. Fields are ubiquitous across a wide range of systems where complex global behaviors can emerge from simple local rules. In particular, we are interested in dynamical and stochastic fields that evolve to achieve specific objectives. We refer to such fields as intelligent fields, reflecting their capacity to model objective-driven behavior in systems ranging from artificial intelligence to neural processes.
@@
Our approach begins with postulating three principles that could underpin such a system. The first principle asserts that there is an evolving configuration that completely characterizes the dynamics of the system. The second principle, locality, suggests that the dynamics occur within a spatial-temporal context. Lastly, the principle of purposefulness proposes that the behavior of the system is directed by its objectives. Although seemingly simple, these principles lead us to develop a framework for objective-driven dynamical stochastic fields, i.e., the intelligent fields. We provide a formal mathematical development of the framework and discuss methodologies for designing the behavior of these fields with a view towards potential applications in artificial intelligence.
@@
The intuition for the first principle relates to the concept of a complete configuration. Consider a system described by a time-varying configuration. If the configuration is complete, meaning that it does not require additional information to determine the system, then the current configuration is sufficient to describe the future. We consider stochastic systems under realism, meaning that these configurations evolve probabilistically and exist independently of observation. Consequently, a complete configuration at any given moment must fully determine the future probabilistic evolution of the system without relying on past configurations, making the dynamics Markovian. In particular, the configuration being complete implies that the transition rule governing the evolution of the configuration can be formulated to be time-invariant, i.e., homogeneous. Otherwise, additional external information would be necessary to define how the transition rules themselves change over time, contradicting the assumption of completeness. The following statement summarizes the above intuition.
@@
It would be quite boring if it were merely a large, plain configuration evolving on its own. Instead, imagine the system to be composed of two entities, each characterized by its own configuration. Each entity’s configuration has two parts: one private and one shared. The private part is accessible only to the entity itself, whereas the shared part is observable by both entities. Thus, at any given moment, each entity observes its own private configuration along with the shared configurations from both entities, but not the other entity’s private configuration. As a result, each entity evolves based solely on this partial information.
@@
Following this intuition, we call this value the objective value, as a generalization of the aforementioned concepts, and we demand our field to satisfy the minimization of this objective value.
@@
In this paper, we present a formal theoretical framework for analyzing and understanding such a system, which we term an intelligent field, reflecting its capacity to model objective-driven behavior. This framework will become concrete in the following sections as we make specific choices to mathematically define the three principles: complete configuration, locality, and purposefulness.
@@
  • Subsequently, in Section 3, we detail the purposefulness principle, complete the dynamical stochastic field with objective-driven behaviors, and explore methodologies for designing the behavior of the system.
@@ 2 Dynamical Stochastic Fields
Let us build the theoretical tools for the dynamical stochastic fields, derived from the first two principles: complete configuration and locality. To begin, we need to provide a concrete definition of the first principle.
@@
Notation. Vectors are represented using bra–ket notation, and operators are denoted in bold font. Additional notation will be introduced throughout the paper as needed.
@@
$\bigl\{\lvert{\omega}\rangle\bigr\}_{{\omega}\in{\Omega}},\qquad\forall{\omega},{\omega}^{\prime}\in{\Omega}:\langle{\omega}\mid{\omega}^{\prime}\rangle=\delta_{{\omega}}^{{\omega}^{\prime}},$ (2.1)
@@
${\mathcal{H}}({\Omega})=\texttt{span}\left(\{\lvert{\omega}\rangle\}_{{\omega}\in{\Omega}}\right).$ (2.2)
@@
The dynamics of the system can be described by dynamics within this Hilbert space. Concretely, elements $\lvert\varphi\rangle$$\in{\mathcal{H}}$ can be used to represent a probabilistic distribution over $\Omega$}.
@@
A vector ${\lvert\varphi\rangle}=\sum_{{\omega}\in{\Omega}}{\varphi}^{{\omega}}\lvert{\omega}\rangle\in{\mathcal{H}}$ is normalized if it represents a probability distribution, i.e.,
$\forall{\omega}:{\varphi}^{{\omega}}\geq 0,\ \quad\text{and}\quad\sum_{{\omega}\in{\Omega}}{\varphi}^{{\omega}}=1.$ (2.3)
Thus, given a normalized vector $\lvert\varphi\rangle$} that describes the probability distribution of a system at a particular time, the probability of finding the system in configuration $\omega$} is
${\varphi}^{{\omega}}=\langle{\omega}\mid\varphi\rangle\in[0,1].$ (2.4)
@@
We can see that the above quantity converges to the delta function $\delta_{{\omega}}^{{\omega}^{\prime}}$ as ${\Delta t}\to 0$. Since the evolution of the system is Markovian, the process can be characterized by its infinitesimal generator. To be self-contained, we define this operator as follows and show that it governs the evolution of the system.
@@
${{\mathbf{G}}}=\sum_{{\omega},{\omega}^{\prime}\in{\Omega}}\lvert{\omega}^{\prime}\rangle G^{{\omega}^{\prime}}_{{\omega}}\langle{\omega}\rvert$, (2.6)
where $G^{{\omega}^{\prime}}_{{\omega}}\in{\mathbb{R}}$. This means that a linear operator $\mathbf{G}$}} is determined by
$G^{{\omega}^{\prime}}_{{\omega}}=\langle{\omega}^{\prime}\rvert{{\mathbf{G}}}\lvert{\omega}\rangle.$ (2.7)
@@
$G^{{\omega}^{\prime}}_{{\omega}}=\langle{\omega}^{\prime}\rvert{{\mathbf{G}}}\lvert{\omega}\rangle=\lim_{{\Delta t}\to 0}\frac{p_{{\Delta t}}({\omega}^{\prime}\mid{\omega})-\delta_{{\omega}}^{{\omega}^{\prime}}}{{\Delta t}}.$ (2.8)
@@
The above infinitesimal generator encapsulates all the information we need to characterize the dynamics of $\lvert\varphi(t)\rangle$, which describes the distribution of system configurations at time $t$. It provides a highly compact formulation of the system dynamics, as shown below. It is called the master equation, or a reformulation of the Kolmogorov forward equations (Kolmogoroff, 1931; Feller, 1940).
@@
$\frac{\mathop{}\!\mathrm{d}}{{\mathop{}\!\mathrm{d}t}}\lvert\varphi(t)\rangle={{\mathbf{G}}}\lvert\varphi(t)\rangle.$ (2.9)
@@
$\lvert\varphi(t)\rangle=e^{{{\mathbf{G}}}t}{\lvert\varphi\rangle}(0).$ (2.10)
Note that the exponential of the linear operator is defined as $e^{{{\mathbf{G}}}t}=1+\sum_{n=1}^{\infty}\frac{1}{n!}({{\mathbf{G}}}t)^{n}$. When the context is clear, we may simply use $\lvert\varphi\rangle$} to denote $\lvert\varphi(t)\rangle$ for notational ease.
Therefore, we have two formulations of the system dynamics: the dynamics of the configuration ${\omega}(t)\in{\Omega}$ itself and the dynamics of the probability distribution $\lvert\varphi(t)\rangle\in{\mathcal{H}}$. To distinguish the two formulations, we adopt the following convention.
Notation. We refer to ${\omega}(t)\in{\Omega}$ as a configuration of the system dynamics and a normalized vector $\lvert\varphi(t)\rangle\in{\mathcal{H}}$ as a state of the system dynamics. A state represents a probability distribution of configurations.
@@ 2.1 A Two-entity View
The locality principle implies that, at time $t$, entity $x$’s immediate next behavior $(\alpha_{t+{\Delta t}},\mu_{t+{\Delta t}})$ depends only on the currently observed local configurations $(\alpha_{t},\beta_{t},\mu_{t})$, and entity $y$’s next behavior $(\beta_{t+{\Delta t}},\nu_{t+{\Delta t}})$ depends only on the currently observed local configurations $(\alpha_{t},\beta_{t},\nu_{t})$.
Formally, the locality principle implies the immediate independence relation, as shown below. Following the notation as stated in equation 2.5, we have
@@
$\langle\alpha^{\prime}\mu^{\prime}\rvert{{\mathbf{M}}}\lvert\alpha\beta\mu\rangle=\lim_{{\Delta t}\to 0}\frac{p_{{\Delta t}}(\alpha^{\prime}\mu^{\prime}\mid\alpha\beta\mu)-\delta_{\alpha\mu}^{\alpha^{\prime}\mu^{\prime}}}{{\Delta t}}.$ (2.15)
@@
$\langle\beta^{\prime}\nu^{\prime}\rvert{{\mathbf{N}}}\lvert\alpha\beta\nu\rangle=\lim_{{\Delta t}\to 0}\frac{p_{{\Delta t}}(\beta^{\prime}\nu^{\prime}\mid\alpha\beta\nu)-\delta_{\beta\nu}^{\beta^{\prime}\nu^{\prime}}}{{\Delta t}}.$ (2.16)
@@
Notation. We adopt the convention that all tensor products are automatically ordered, e.g., ${\mathcal{H}}({\mathcal{A}})\otimes{\mathcal{H}}({\mathcal{M}})\otimes{\mathcal{H}}({\mathcal{B}})\otimes{\mathcal{H}}({\mathcal{N}})$ is equivalent to ${\mathcal{H}}({\mathcal{A}})\otimes{\mathcal{H}}({\mathcal{B}})\otimes{\mathcal{H}}({\mathcal{M}})\otimes{\mathcal{H}}({\mathcal{N}})$. Therefore, the tensor product of basis vectors $\lvert\alpha\mu\rangle\in{\mathcal{H}}({\mathcal{A}}\times{\mathcal{M}})$ and $\lvert\beta\nu\rangle\in{\mathcal{H}}({\mathcal{B}}\times{\mathcal{N}})$ is $\lvert\alpha\mu\rangle\otimes\lvert\beta\nu\rangle=\lvert\alpha\beta\mu\nu\rangle\in{\mathcal{H}}$. Formally, this means we work in a quotient space obtained by identifying tensor products differing only by ordering. For ease of notation, however, we omit explicit references to this quotient operation and simply adopt this ordering as a standard convention throughout this paper.
@@
Given infinitesimal generator ${{\mathbf{M}}}\colon{\mathcal{H}}({\mathcal{A}}\times{\mathcal{B}}\times{\mathcal{M}})\to{\mathcal{H}}({\mathcal{A}}\times{\mathcal{M}})$ of the entity $x$, and ${{\mathbf{N}}}\colon{\mathcal{H}}({\mathcal{A}}\times{\mathcal{B}}\times{\mathcal{N}})\to{\mathcal{H}}({\mathcal{B}}\times{\mathcal{N}})$ of the entity $y$, we define their embeddings $\bar{{\mathbf{M}}},\bar{{\mathbf{N}}}$ in the space ${\mathcal{B}}({\mathcal{H}},{\mathcal{H}})$ as the following. Given basis vectors $\lvert\alpha\beta\mu\nu\rangle\in{\mathcal{H}}$,
$\bar{{\mathbf{M}}}\lvert\alpha\beta\mu\nu\rangle$ $:=({{\mathbf{M}}}\lvert\alpha\beta\mu\rangle)\otimes\lvert\beta\nu\rangle,$ (2.18)
$\bar{{\mathbf{N}}}\lvert\alpha\beta\mu\nu\rangle$ $:=({{\mathbf{N}}}\lvert\alpha\beta\nu\rangle)\otimes\lvert\alpha\mu\rangle.$ (2.19)
@@
We have so far considered the case of a system consisting of two entities. Next, we extend this theoretical model to multiple interacting entities that collectively form a field.